OpenStudy (anonymous):

Describe the transformation for : y= -x²+6x-5

OpenStudy (anonymous):

The parent function is the squaring function, \(y=x^2\) Look at all the ways \(y= -x^2+6x-5\) is different from the parent function. Putting it in vertex form will also help.

OpenStudy (anonymous):

how do i do that ? :$

OpenStudy (anonymous):

Vertex form? Complete the square.

OpenStudy (anonymous):

ohhh okay !

OpenStudy (anonymous):

\[\huge -(x-3)^2+4\]

OpenStudy (anonymous):

That looks right.

OpenStudy (anonymous):

\[\huge Range = {\epsilon \mathbb{R}| y \neq 0}\]

OpenStudy (anonymous):

Not for the function you have above, no.

OpenStudy (anonymous):

y doesnot equal 4** sorry typo

OpenStudy (anonymous):

y can equal 4 here, just not anything above it.

OpenStudy (anonymous):

my solution says it says y can equal any real number :S

OpenStudy (anonymous):

yeah i thought it was that too,

OpenStudy (anonymous):

Not for a parabola it can't.

OpenStudy (anonymous):

The domain is any real number for x...

OpenStudy (anonymous):

yeah that's what i thought, look at number 8 http://mail.wecdsb.on.ca/~janet_maitland/FOV1-0006CFBC/FOV1-0006E9C7/unit%20three%20practice%20answers.pdf?Plugin=Loft

OpenStudy (anonymous):

maybe she made a mistake :S

OpenStudy (anonymous):

Yeah, you can see clearly from the graph that the range is y≤4.

OpenStudy (anonymous):

okay thanks :D i thought i might have been doing to wrong